A. Illustrate the equations that describe the physical quantities of motion:


distance, displacement, velocity and acceleration.

1
Expert's answer
2021-11-11T03:33:01-0500

1)V = U + at

2)V2= U2 + 2aS

2)S = Ut + 1/2(at2)

Where

S=distance/displacement

V&U=Initial and final velocity respectively

a=acceleration